# Restriction Estimates for 2D Surfaces of Finite Type 3 and Applications to Dispersive Equations

### v2

@@ Abstract.
In this paper, we prove the restriction estimates for 2D surfaces $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}\pm\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ by reducing to Wang–Wu’s result on the perturbed paraboloid and Demeter–Wu’s result on the perturbed hyperboloid. The method is based on the rescaling technique developed in [LMZ21]. Besides, we will use the estimates to give a better analysis for discrete nonlinear Schrödinger equations.
@@ 1. Introduction
by assuming Conjecture 1.1. Thus, it is of great interest to study the “hyperbolic” case. However, this case is more difficult than the previous one, as the two principal curvatures have different signs. To overcome this difficulty, we invoke a “hyperbolic” analogue of Theorem 1.3, established by Demeter and Wu in [DW25].
@@
$\|E_{S}g\|_{L^{p}(\mathbb{R}^{3})}\leq C\|g\|_{L^{p}({[0,1]^{2}})}.$
@@
Let $\varphi_{1},\varphi_{2}$ be of finite type $3$, then for $S:=\{(\xi_{1},\xi_{2},\varphi_{1}(\xi_{1})-\varphi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, we have the local restriction estimate
@@
Naturally, the prototypical example this time is $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}-\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$. It is also worth mentioning that we can only obtain the estimates for surfaces such as $\{(\xi_{1},\xi_{2},\xi_{1}^{m_{1}}+\xi_{2}^{m_{2}}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ from [BMV17]. Therefore, at present, we have the Tomas–Stein type estimate (the point H in Figure 2) and the trivial $L^{1}\to L^{\infty}$ bound (the point D). Now, by interpolation, the range of $(\frac{1}{q},\frac{1}{p})$ shrinks to the quadrilateral ACDH (excluding the segments $\overline{AH}$). The Tomas–Stein type estimate can be derived from oscillatory integrals, which were extensively studied in [IM11, IM16, IKM10].
@@
$\|Eg\|_{L^{22/7}(B_{R}^{3})}\leq Q(R)\|g\|_{L^{\infty}([0,1]^{2})},\quad\forall g\in L^{\infty}([0,1]^{2}).$ (1.2)
@@ 2. Restriction estimate on $\Omega_{\lambda}$
$\left\|Eg_{\Omega_{\lambda}}\right\|_{L^{22/7}(B_{R}^{3})}\lesssim_{\varepsilon}\left(\sum_{\tau}\left\|Eg_{\tau\cap\Omega_{\lambda}}\right\|_{L^{22/7}(B_{R}^{3})}^{2}\right)^{1/2}.$ (2.1)
@@
$\left\{\begin{aligned} &\widetilde{x_{1}}:=\lambda^{-1/2}K^{-1/2}x_{1}+3\lambda^{3/2}K^{-1/2}x_{3},\;\widetilde{x_{2}}:=K^{-1/3}x_{2},\;\widetilde{x_{3}}:=K^{-1}x_{3},\\ &\widetilde{g}(\eta_{1},\eta_{2}):=\lambda^{-1/2}K^{-5/6}g(\lambda+\lambda^{-1/2}K^{-1/2}\eta_{1},K^{-1/3}\eta_{2}),\\ &\Gamma(\eta_{1},\eta_{2}):=3\eta_{1}^{2}+\eta_{2}^{3}+\lambda^{-3/2}K^{-1/2}\eta_{1}^{3}.\end{aligned}\right$. (2.3)
@@
A key observation is that we can further decompose $0,1]^{2}$ into two parts, one away from $\eta_{2}=0$ and one near $\eta_{2}=0$ (see Figure 5). Then, we can obtain an estimate for $E_{[0,1]^{2}}^{\Gamma}$ as follows. The idea is inspired by [LZ23].
@@
$\|E_{S}\widetilde{f}_{\widetilde{\Omega_{0}}}\|_{L^{22/7}(B_{R}^{3})}\leq A_{0}(\varepsilon)R^{\varepsilon}\|\widetilde{f}\|_{L^{\infty}([0,1]^{2})},\quad\forall\widetilde{f}\in L^{\infty}([0,1]^{2}).$ (2.4)
@@
$\|E_{S}\widetilde{f}_{\widetilde{\Omega_{1}}}\|_{L^{22/7}(B_{R}^{3})}\lesssim_{\delta}K^{\delta}\left(\sum_{\widetilde{\tau}}\|E_{S}\widetilde{f}_{\widetilde{\tau}}\|_{L^{22/7}(B_{R}^{3})}^{2}\right)^{1/2},\quad\forall\delta>0,$ (2.5)
@@
$\|E_{S}\widetilde{f}_{\widetilde{\tau}}\|_{L^{22/7}(B_{R}^{3})}\lesssim K^{7/12}\|\widetilde{E}_{[0,1]^{2}}^{\overline{\Gamma}}\overline{f}\|_{L^{p}(B_{RK^{-1/3}}^{3})}\lesssim K^{-1/4}A\left(\frac{R}{K^{1/3}}\right)\|\widetilde{f}\|_{L^{\infty}([0,1]^{2})}.$
@@
The proof for the surface $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}+\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ also applies to the surface $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}-\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$.
@@
  • Let $\varphi_{1}$ be of finite type $2$ and $\varphi_{2}$ be of finite type $3$, then for $S:=\{(\xi_{1},\xi_{2},\phi_{1}(\xi_{1})\pm\phi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, we have the local restriction estimate
@@
  • For $S:=\{(\xi_{1},\xi_{2},\phi_{1}(\xi_{1})-\phi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, we have the Tomas–Stein type estimate (the point H in Figure 8) instead. Then, still by interpolation, we can prove the case when ($\frac{1}{q},\frac{1}{p}$) lies in the quadrilateral ACDH (excluding the segment $\overline{AH}$).
@@ 3. Applications to dispersive equations
The discrete nonlinear equations (or, to be specific, nonlinear equations on lattices) have served as models for many important physical phenomena. For example, the non–exponential energy relaxation in solids [CP89] and the self–focusing and collapse of Langmuir waves in plasmas [Zak83]. For further physical interest, we refer to [KRB01, Kev09]. On the other hand, DNLS also has its own mathematical interest. In fact, one often discretizes PDEs for computer simulations or approximations. Then, establishing the theory of the continuum limit, i.e., whether the solutions of the discrete equations converge to the solutions of the original equations, is very essential. For more recent progress on continuum limit theories, see [KLS13, HY19, CA23, Wan25].
@@
$\left\{\begin{aligned} &i\partial_{t}u(x,t)+\Delta_{disc}u(x,t)=\mu|u|^{\alpha-1}u,\\ &u(x,0)=f(x),\quad(x,t)\in{\mathbb{Z}}^{d}\times{\mathbb{R}},\end{aligned}\right$. (3.1)
@@
$\Delta_{disc}u=\mathcal{F}_{disc}^{-1}\left\{\left[\sum_{i=1}^{d}4\sin^{2}(\frac{\xi_{i}}{2})\right]\mathcal{F}_{disc}u\right\}\triangleq\mathcal{F}_{disc}^{-1}\left\{\omega(\xi)\mathcal{F}_{disc}u\right\}.$ (3.2)
@@
$\left\{\begin{aligned} &i\partial_{t}u(x,t)+\Delta u(x,t)=F(x,t),\\ &u(x,0)=f(x),\quad(x,t)\in{\mathbb{R}}^{d}\times{\mathbb{R}}.\end{aligned}\right$. (3.3)
@@
$\|e^{it\Delta_{disc}}f\|_{L_{t}^{q}\ell_{x}^{r}(\mathbb{R}\times{\mathbb{Z}}^{d})}\lesssim\|f\|_{\ell^{2}({\mathbb{Z}}^{d})},$ (3.4)
@@
$\|\int_{s<t}e^{i(t-s)\Delta_{disc}}F(\cdot,s)ds\|_{L_{t}^{q}\ell_{x}^{r}(\mathbb{R}\times{\mathbb{Z}}^{d})}\lesssim\|F\|_{L_{t}^{\widetilde{q}^{\prime}}\ell_{x}^{\widetilde{r}^{\prime}}(\mathbb{R}\times{\mathbb{Z}}^{d})},$ (3.5)
@@
$\|e^{it\Delta_{disc}}f\|_{L^{p}\left(I;\ell^{p}\right)}\leq\eta,$ (3.6)
@@
$\left\{\begin{aligned} &\sup_{t\in[0,T^{\ast})}\|u(t)\|_{\widehat{L^{q^{\prime}}}}=\infty,\quad if\;\alpha<p\leq\alpha q^{\prime},\\ &\|u\|_{L^{p}\left([0,T^{\ast});\ell^{p}\right)}=\infty,\quad if\;\alpha=p.\end{aligned}\right$.
@@
$\left\{\begin{aligned} &0\leq\frac{1}{\alpha}<-\frac{13}{55}\cdot\frac{1}{q}+\frac{7}{22},\quad if\;\;0\leq\frac{1}{q}<\frac{1}{2},\\ &0\leq\frac{1}{\alpha}\leq\frac{1}{5},\quad if\;\;\frac{1}{q}=\frac{1}{2},\end{aligned}\right$.

### v3

@@ Abstract.
In this paper, we prove the restriction estimates for 2D surfaces $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}\pm\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ by reducing to Wang–Wu’s result on the perturbed paraboloid and to the results on the perturbed hyperboloid obtained by Buschenhenke, Müller, and Vargas, as well as by Guo and Oh. The method is based on the rescaling technique developed in [LMZ21]. Besides, we will use the estimates to give a better analysis for discrete nonlinear Schrödinger equations.
@@ 1. Introduction
by assuming Conjecture 1.1. Thus, it is of great interest to study the “hyperbolic” case. However, this case is more difficult than the previous one, as the two principal curvatures have different signs. To overcome this difficulty, we invoke a hyperbolic analogue of Theorem 1.3, established independently by Buschenhenke, Müller, and Vargas, and by Guo and Oh (see [BMV23, GO24]).
@@
then for $p>13/4$, $p>2q^{\prime}$, we have a restriction estimate, for $C=C(p)>0$ independent of $\phi$,
$\|E_{S}g\|_{L^{p}(\mathbb{R}^{3})}\leq C\|g\|_{L^{q}({[0,1]^{2}})}.$
@@
Let $\varphi_{1},\varphi_{2}$ be of finite type $3$, then for $S:=\{(\xi_{1},\xi_{2},\varphi_{1}(\xi_{1})-\varphi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, $\frac{3}{10}\leq\frac{1}{p}<\frac{4}{13}$, $\frac{11}{2p}+\frac{3}{4}\leq\frac{5}{2q^{\prime}}$, we have the local restriction estimate
$\left\|E_{S}g\right\|_{L^{p}(B_{R}^{3})}\lesssim_{\varepsilon}R^{\varepsilon}\|g\|_{L^{q}([0,1]^{2})},\quad\forall\varepsilon>0.$
@@
Naturally, the prototypical example this time is $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}-\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$. It is also worth mentioning that we can only obtain the estimates for surfaces such as $\{(\xi_{1},\xi_{2},\xi_{1}^{m_{1}}+\xi_{2}^{m_{2}}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ from [BMV17]. Therefore, at present, we have the Tomas–Stein type estimate (the point H in Figure 2) and the trivial $L^{1}\to L^{\infty}$ bound (the point D). Now, by interpolation, the range of $(\frac{1}{q},\frac{1}{p})$ shrinks to the pentagon ACDHK (excluding the segments $\overline{AK}$, $\overline{KH}$). The Tomas–Stein type estimate can be derived from oscillatory integrals, which were extensively studied in [IM11, IM16, IKM10].
@@
(1.2) $\|Eg\|_{L^{22/7}(B_{R}^{3})}\leq Q(R)\|g\|_{L^{\infty}([0,1]^{2})},\quad\forall g\in L^{\infty}([0,1]^{2}).$
@@ 2. Restriction estimate on $\Omega_{\lambda}$
(2.1) $\left\|Eg_{\Omega_{\lambda}}\right\|_{L^{22/7}(B_{R}^{3})}\lesssim_{\varepsilon}\left(\sum_{\tau}\left\|Eg_{\tau\cap\Omega_{\lambda}}\right\|_{L^{22/7}(B_{R}^{3})}^{2}\right)^{1/2}.$
@@
(2.3) $\left\{\begin{aligned} &\widetilde{x_{1}}:=\lambda^{-1/2}K^{-1/2}x_{1}+3\lambda^{3/2}K^{-1/2}x_{3},\;\widetilde{x_{2}}:=K^{-1/3}x_{2},\;\widetilde{x_{3}}:=K^{-1}x_{3},\\ &\widetilde{g}(\eta_{1},\eta_{2}):=\lambda^{-1/2}K^{-5/6}g(\lambda+\lambda^{-1/2}K^{-1/2}\eta_{1},K^{-1/3}\eta_{2}),\\ &\Gamma(\eta_{1},\eta_{2}):=3\eta_{1}^{2}+\eta_{2}^{3}+\lambda^{-3/2}K^{-1/2}\eta_{1}^{3}.\end{aligned}\right$.
@@
A key observation is that we can further decompose $0,1]^{2}$ into two parts, one away from $\eta_{2}=0$ and one near $\eta_{2}=0$ (see Figure 5). Then, we can obtain an estimate for $E_{[0,1]^{2}}^{\Gamma}$ as follows. The idea is inspired by [LZ26].
@@
(2.4) $\|E_{S}\widetilde{f}_{\widetilde{\Omega_{0}}}\|_{L^{22/7}(B_{R}^{3})}\leq A_{0}(\varepsilon)R^{\varepsilon}\|\widetilde{f}\|_{L^{\infty}([0,1]^{2})},\quad\forall\widetilde{f}\in L^{\infty}([0,1]^{2}).$
@@
(2.5) $\|E_{S}\widetilde{f}_{\widetilde{\Omega_{1}}}\|_{L^{22/7}(B_{R}^{3})}\lesssim_{\delta}K^{\delta}\left(\sum_{\widetilde{\tau}}\|E_{S}\widetilde{f}_{\widetilde{\tau}}\|_{L^{22/7}(B_{R}^{3})}^{2}\right)^{1/2},\quad\forall\delta>0,$
@@
$\|E_{S}\widetilde{f}_{\widetilde{\tau}}\|_{L^{22/7}(B_{R}^{3})}\lesssim K^{7/12}\|\widetilde{E}_{[0,1]^{2}}^{\overline{\Gamma}}\overline{f}\|_{L^{22/7}(B_{RK^{-1/3}}^{3})}\lesssim K^{-1/4}A\left(\frac{R}{K^{1/3}}\right)\|\widetilde{f}\|_{L^{\infty}([0,1]^{2})}.$
@@
The proof for the surface $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}+\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$ also applies to $S:=\{(\xi_{1},\xi_{2},\xi_{1}^{3}-\xi_{2}^{3}):(\xi_{1},\xi_{2})\in[0,1]^{2}\}.$ Moreover, the condition $\frac{11}{2p}+\frac{3}{4}\leq\frac{5}{2q^{\prime}}$ arises naturally from the same calculation after replacing $22/7$ and $\infty$ with $p$ and $q$, respectively.
@@
  • Let $\varphi_{1}$ be of finite type $2$ and $\varphi_{2}$ be of finite type $3$, then for $S:=\{(\xi_{1},\xi_{2},\phi_{1}(\xi_{1})+\phi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, we have the local restriction estimate
@@
    In the same spirit, for $S:=\{(\xi_{1},\xi_{2},\phi_{1}(\xi_{1})-\phi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, $\frac{3}{10}\leq\frac{1}{p}<\frac{4}{13}$, $\frac{11}{2p}+\frac{3}{4}\leq\frac{5}{2q^{\prime}}$, we have the local restriction estimate
    $\left\|E_{S}g\right\|_{L^{p}(B_{R}^{3})}\lesssim_{\varepsilon}R^{\varepsilon}\|g\|_{L^{q}([0,1]^{2})},\quad\forall\varepsilon>0.$
@@
  • For $S:=\{(\xi_{1},\xi_{2},\phi_{1}(\xi_{1})-\phi_{2}(\xi_{2})):(\xi_{1},\xi_{2})\in[0,1]^{2}\}$, we have the Tomas–Stein type estimate (the point H in Figure 8) instead. Then, still by interpolation, we can prove the case when ($\frac{1}{q},\frac{1}{p}$) lies in the pentagon ACDHK (excluding the segments $\overline{AK}$, $\overline{KH}$).
@@ 3. Applications to dispersive equations
The discrete nonlinear equations (or, to be specific, nonlinear equations on lattices) have served as models for many important physical phenomena. For example, the non–exponential energy relaxation in solids [CP89] and the self–focusing and collapse of Langmuir waves in plasmas [Zak83]. For further physical interest, we refer to [KRB01, Kev09]. On the other hand, DNLS also has its own mathematical interest. In fact, one often discretizes PDEs for computer simulations or approximations. Then, establishing the theory of the continuum limit, i.e., whether the solutions of the discrete equations converge to the solutions of the original equations, is very essential. For more recent progress on continuum limit theories, see [KLS13, HY19, CA23, Wan26].
@@
(3.1) $\left\{\begin{aligned} &i\partial_{t}u(x,t)+\Delta_{disc}u(x,t)=\mu|u|^{\alpha-1}u,\\ &u(x,0)=f(x),\quad(x,t)\in{\mathbb{Z}}^{d}\times{\mathbb{R}},\end{aligned}\right$.
@@
(3.2) $\Delta_{disc}u=\mathcal{F}_{disc}^{-1}\left\{\left[\sum_{i=1}^{d}4\sin^{2}(\frac{\xi_{i}}{2})\right]\mathcal{F}_{disc}u\right\}\triangleq\mathcal{F}_{disc}^{-1}\left\{\omega(\xi)\mathcal{F}_{disc}u\right\}.$
@@
(3.3) $\left\{\begin{aligned} &i\partial_{t}u(x,t)+\Delta u(x,t)=F(x,t),\\ &u(x,0)=f(x),\quad(x,t)\in{\mathbb{R}}^{d}\times{\mathbb{R}}.\end{aligned}\right$.
@@
(3.4) $\|e^{it\Delta_{disc}}f\|_{L_{t}^{q}\ell_{x}^{r}(\mathbb{R}\times{\mathbb{Z}}^{d})}\lesssim\|f\|_{\ell^{2}({\mathbb{Z}}^{d})},$
@@
(3.5) $\|\int_{s<t}e^{i(t-s)\Delta_{disc}}F(\cdot,s)ds\|_{L_{t}^{q}\ell_{x}^{r}(\mathbb{R}\times{\mathbb{Z}}^{d})}\lesssim\|F\|_{L_{t}^{\widetilde{q}^{\prime}}\ell_{x}^{\widetilde{r}^{\prime}}(\mathbb{R}\times{\mathbb{Z}}^{d})},$
@@
(3.6) $\|e^{it\Delta_{disc}}f\|_{L^{p}\left(I;\ell^{p}\right)}\leq\eta,$
@@
$\left\{\begin{aligned} &\sup_{t\in[0,T^{\ast})}\|u(t)\|_{\widehat{L^{q^{\prime}}}}=\infty,\quad\text{if}\;\alpha<p\leq\alpha q^{\prime},\\ &\|u\|_{L^{p}\left([0,T^{\ast});\ell^{p}\right)}=\infty,\quad\text{if}\;\alpha=p.\end{aligned}\right$.
@@
$\left\{\begin{aligned} &0<\frac{1}{\alpha}<\frac{4}{13},\quad\text{if}\;\;0<\frac{1}{q}\leq\frac{3}{130},\\ &0<\frac{1}{\alpha}<\frac{4}{13}-\frac{7}{31}\left(\frac{1}{q}-\frac{3}{130}\right),\quad\text{if}\;\;\frac{3}{130}<\frac{1}{q}\leq\frac{1}{2},\\ &0<\frac{1}{\alpha}\leq\frac{1}{5},\quad\text{if}\;\;q=2,\end{aligned}\right$.